OpenStudy (anonymous):

Where AH bisects angleMAT , MangleMAH=3(x-4) and MangleHAT= 1/2(4x+6). Find x Find MangleHAT Find MangleMAT

jhonyy9 (jhonyy9):

can you drawing it please ?

jhonyy9 (jhonyy9):

these are angles inside a triangle ?

jhonyy9 (jhonyy9):

so than i understand you right MAT is an angle what is bisected by AH so than angle MAH has measure =3(x-4) and the measure of HAT = 1/2(4x+6) - so do you know what mean AH bisect angle MAT ?

jhonyy9 (jhonyy9):

so because AH is bisector so this mean that angle MAH and angle HAT has equal measures - yes ? so can you continue it now ?

jhonyy9 (jhonyy9):

so this mean that 3(x-4) = 1/2(4x+6) from this can you calculi the x ?

OpenStudy (anonymous):

its 1/2(4x+6)

jhonyy9 (jhonyy9):

there is something wrong - so the measure of angle HAT how is it correct ? (4x+6)/2 or 1/(2*(4x+6)) ?

jhonyy9 (jhonyy9):

so this mean (4x+6) divided by 2 - yes ?

OpenStudy (anonymous):

i really dont know sir

jhonyy9 (jhonyy9):

in your book has in this form ?

OpenStudy (anonymous):

no

jhonyy9 (jhonyy9):

2x+3=3x-12 15=x sorry above i have missed one minus in plus so x=15 will be right than x=15 measure angle MAH = 3(x-4) = 3x-12 = 3*15-12 = 45-12 = 33 can you calculi now the measure of angle HAT ?

jhonyy9 (jhonyy9):

so you know that (4x+6) divided by 2 = 2x+3 so than just substitute this value of x, x=15, in place of x and what will get ?

jhonyy9 (jhonyy9):

courage please - this is easy

jhonyy9 (jhonyy9):

m<HAT = 2*15+3 = ?

jhonyy9 (jhonyy9):

so m<HAT = m<MAH what i have calculated above first and have got 33 because 2*15+3=30+3=33 yes ?

jhonyy9 (jhonyy9):

hope so much that you understand it now sure right - yes ?

OpenStudy (anonymous):

ok i understand